# Setting Limits on Blazar-Boosted Dark Matter with Xenon-based Detectors

### v1

@@ Abstract
Dual-phase xenon time projection chambers achieve optimal sensitivity for dark matter in the 10 to $1000\text{\,}\mathrm{G}\mathrm{eV}\mathrm{c}^{-2}$ mass range, but sub-GeV DM particles lack sufficient energy to produce nuclear recoils above detection thresholds in these detectors. Blazar-boosted dark matter offers a way to overcome this limitation. Relativistic jets in active galactic nuclei can accelerate light dark matter in their host-galaxy halos to energies that can leave detectable nuclear recoil signals in xenon-based detectors on Earth. We present the first blazar-boosted dark matter search that incorporates detector response modeling, using public data from XENON1T and LZ for the blazar TXS 0506+056. We model dark matter-proton scattering in the jet environment, covering the full process from jet acceleration through to detector response. We explore how the host galaxy dark matter density profile impacts our analysis. We exclude dark matter–nucleon cross-sections for  $1\text{\,}\mathrm{MeV}$ dark matter between $5.8\text{\times}{10}^{-31}\text{\,}{\mathrm{cm}}^{2}$ and $6.3\text{\times}{10}^{-29}\text{\,}{\mathrm{cm}}^{2}$ using XENON1T data and between $9.9\text{\times}{10}^{-32}\text{\,}{\mathrm{cm}}^{2}$ and $2.5\text{\times}{10}^{-28}\text{\,}{\mathrm{cm}}^{2}$ from LZ EFT searches.
Our results show that astrophysical uncertainties, especially the dark matter density profile near the supermassive black hole, set the primary limitation on the sensitivity rather than detector systematics. These findings highlight blazar-boosted dark matter as a promising probe of light dark matter and motivate further astrophysical modeling.
@@ I Introduction
The dark matter (DM) problem emerged from astronomical and cosmological observations that reveal a discrepancy between the observed gravitational effects and the visible matter in the universe. This discrepancy requires either additional undetected mass or modifications to our current understanding of gravity. One proposed explanation involves dark matter particles, a form of non-baryonic matter [1].
@@
Active Galactic Nuclei (AGNs) are the luminous central regions of galaxies powered by supermassive black holes. Blazars represent a special subclass of AGNs where jets are oriented toward Earth, making them appear exceptionally bright [9, 10].
@@
This paper is organized as follows: in Sec. II, we present the selection criteria for the blazar as a boosted-DM source and model both its jet physics and the DM profile of the host galaxy. In Sec. III, we calculate the BBDM flux, including the DM upscattering mechanism within the jet and Earth attenuation effects. Sec. IV describes the nuclear recoil interactions in liquid xenon detectors, essentially computing the recoil spectra. In Sec. V, we detail our approach for setting constraints on blazar-boosted dark matter using publicly available data from the XENON1T and LZ detectors. Sec. VI presents our conclusions.
@@ II.1 Selection of blazar-boosting source
In 2013, the IceCube Neutrino Observatory [11] detected the first high-energy astrophysical neutrino with energies exceeding 1 $\text{\,}\mathrm{P}\mathrm{e}\mathrm{V} [1$2]. Since then, it has compiled a growing list of such events (see Ref. [13] and references therein). Yet, so far, only the Galactic plane [14] (at 4.5 $\sigma$), and two astronomical objects have been associated with high-energy neutrinos detected by IceCube with a significance greater than $\sim 3\,\sigma$: the blazar TXS 0506+056  at $z=0.3365$ (at the 3–3.5 $\sigma$ level  [15, 16]) and the local Seyfert 2 galaxy NGC 1068 at $z=0.004$ (at the $4.2\,\sigma$ level  [17, 18]).
@@
Since light travel time from this blazar is 3.75 Gyr [19], any particles it emits must travel at least this long to reach Earth. Given that the Universe is 13.8 Gyr old, and that galaxies might form 0.2 Gyr after the Big Bang [e.g., 20], the blazar’s maximum possible age is about 9.8 Gyr (13.83.750.2=9.813.8-3.75-0.2=9.8 Gyr). For BBDM particles to be detectable on Earth, they must complete their journey within this 9.8 Gyr timeframe. This constraint establishes a minimum velocity requirement for these particles, corresponding to a Lorentz factor of at least 1.1. Since this is very close to $\gamma=1$ (i.e., the non-relativistic limit), this condition effectively places no meaningful constraint on the particles’ velocities.
@@ II.2 Blazar jet physics
The redshift of the source, and consequently its distance, has been firmly determined by Ref. [19].
@@
The mass of the central black hole (BH) in TXS 0506+056  is estimated to be $M_{\mathrm{BH}}\approx 6.3\times 10^{8}M_{\odot}$, where $M_{\odot}$ denotes the solar mass. This estimate is based on the assumption that the host galaxy is a typical giant elliptical with an absolute R magnitude of $M(R)\approx-22.9$ [21]. Consequently, this corresponds to an Eddington power $L_{\mathrm{Edd}}\approx 7.9\times 10^{46}$ erg s-1.
@@
The motion of the relativistic particles in the blazar jet can be approximately described according to the “blob geometry” [22] in which the accelerated particles are confined within spherical regions (blobs) that move with speed $v_{{\mathrm{B}}}=\beta_{{\mathrm{B}}}c$, and bulk Lorentz factor $\Gamma_{{\mathrm{B}}}=(1-\beta_{{\mathrm{B}}}^{2})^{-1/2}$ along the jet axis. The latter is, in general, inclined at an angle $\theta_{\text{l.o.s.}}$ with respect to our line-of-sight (l.o.s.). Within each blob, the individual particles move isotropically relative to the blob center of mass frame.
@@
In our analysis, we adopt the lepto-hadronic model used in [23, 24] to describe the jet emission of blazar TXS 0506+056. This model characterizes the electromagnetic energy spectrum, i.e., the energy of the photons emitted by the accelerated charged particles (primarily electrons and protons) in the blazar’s jet. The model assumes a homogeneous and isotropic proton energy spectrum, following a single power-law with slope $\alpha_{p}$ within the energy range $\gamma^{\prime}_{{\mathrm{min}},p}\leq E_{p}^{\prime}/m_{p}\leq\gamma^{\prime}_{{\mathrm{max}},p}$, where $E^{\prime}_{p}$ is the proton energy in the blob’s rest frame. By fitting the observed electromagnetic spectral energy distribution, Ref. [23, 24] determines the bulk Lorentz factor $\Gamma_{\rm B}$, the spectral index $\alpha_{p}$, the minimum and maximum Lorentz factors $\gamma^{\prime}_{{\mathrm{min}},p}$ and $\gamma^{\prime}_{{\mathrm{max}},p}$, and the proton luminosity $L_{p}$ that normalizes the energy power-law spectrum.
@@
The angle $\theta_{\text{l.o.s.}}$ for blazars is generally known to be small, with $\theta_{\text{l.o.s.}}<15-20^{\circ}$ (e.g. Ref. [25, 10]). However, precise values for individual objects are uncertain and difficult to derive. In the analysis of [23, 24], it is assumed that the bulk Lorentz factor of the jet $\Gamma_{\rm B}$ equals $\mathcal{D}/2$, where the latter is the Doppler factor $\mathcal{D}\equiv[\Gamma_{\rm B}\left(1-\beta_{\rm B}\cos\theta_{\text{l.o.s.}}\right)]^{-1}$. This implies $\theta_{\text{l.o.s.}}=0^{\circ}$, which is highly improbable and would preclude superluminal motion in this source [26]. This contradicts observations, as superluminal motion with $\beta_{\rm app}=1.07\pm 0.14$ has been measured [27]. A further complication is that the jet parameters are very likely frequency-dependent, with $\Gamma_{\rm B}$ and $\mathcal{D}$ being much larger at high energies compared to the radio band (e.g. [26]).
@@
We therefore assume $\mathcal{D}=35$, which is the smallest value for this parameter derived by the lepto-hadronic spectral energy distribution modeling of [24], as this implies the largest possible $\theta_{\text{l.o.s.}}=\arcsin(1/\mathcal{D})\sim 1.64^{\circ}$ (see Appendix A of [25]). For this choice, and based on the definition $\mathcal{D}$, it can be shown that $\Gamma_{\rm B}=\mathcal{D}$. This means that the lowest value for the proton power in Tab. 1 of Ref. [24], which has been derived assuming $\Gamma_{\rm B}=\mathcal{D}/2$, needs to be multiplied by a factor of 4 for consistency, giving $L_{p}=6.4\times 10^{48}\,\text{erg}/\text{s}$. We take the other blazar jet parameters at the best-fit values given in [24]. The entire list of parameters and corresponding values considered in our work are summarized in Tab. 1.
@@
With the parameters in the table, we estimate the number of protons ejected per time, kinetic energy $T_{p}$ and solid angle $\Omega_{p}$, in the rest frame of an observer that sees the blobs moving with velocity $\beta_{{\mathrm{B}}}$ along the jet axis, as [7]
@@
where $m_{p}\simeq 0.938\,\text{GeV}/c^{2}$ is the proton mass, $\beta_{p}=[1-m^{2}_{p}/(T_{p}+m_{p})^{2}]^{1/2}$ is the proton speed in units of $c$, $\mu$ is the cosine of the angle of motion with respect to the jet axis, and $c_{p}$ is a normalization constant, which we fix by requiring the matching with the fitted proton luminosity $L_{p}$ [8].
@@
Table 1: Parameters currently used in the BBDM code for TXS 0506+056 [23, 24]. These parameters were partially derived from [7, 8]. The quantities flagged with a star () correspond to mean values computed from the ranges given in the second column of Table 1 of Ref. [24]. The resulting values of the normalization constants $c_{e,\,p}$, as well as the redshift $z$ [28], luminosity distance $d_{L}$ for $H_{0}=70$, $\Omega_{\rm m}=0.3$, and $\Omega_{\rm\Lambda}=0.7$, BH mass $M_{{\mathrm{BH}}}$ [21].
@@ II.3 Dark Matter Density profile
Navarro–Frenk–White. The NFW profile [29] is often used as the default model for galactic DM halos. It takes the form:
@@
where $\rho_{s}$ is the characteristic density (a normalization parameter that sets the density scale) and and $r_{s}$ is the characteristic scale radius (which determines the transition point between the inner and outer regions of the density profile). The characteristic density $\rho_{s}$ is defined [30] by requiring that the total mass within the virial radius $r_{\text{vir}}$ equals the virial mass:
@@
where $\delta_{\text{vir}}\simeq 100$ is the virial overdensity in the standard cosmological model [31], and $\rho_{c}=1.053672\times 10^{-5}\,h^{-2}\,(\text{GeV}/c^{2})\,\text{cm}^{-3}$ is the critical density of the Universe. Here, $h\simeq 0.674$ is the Hubble expansion rate in units of 100 km/s/Mpc [32].
@@
The characteristic radius $r_{s}$ and the virial radius $r_{\text{vir}}$ entering Eq.  (5) are expressed in terms of the virial mass as[33, 34]
@@
The virial mass of the DM halo is estimated via the following relation which stems from numerical simulations [35] (see also [36, 37]):
@@
giving $M_{\text{vir}}\simeq 5.2\times 10^{12}M_{\odot}$ for the parameters in Tab. 1. Using this value in Eqs. (7) and (6), we get $r_{\text{vir}}\simeq 273\,\text{kpc}$ and $r_{s}\simeq 54\,\text{kpc}$, while from Eq. (5), $\rho_{s}\simeq 0.10\,(\text{GeV}/c^{2})\,\text{cm}^{-3}$.
The strength of the NFW profile lies in its close agreement with observed rotation velocity curves at large radii. However, at small radii, the DM profile is poorly constrained, making the NFW profile an extrapolation of large-scale trends to small scales without strong empirical support. Furthermore, numerical simulations indicate that the NFW profile is not expected to hold at small radii, since baryonic processes such as stellar feedback and gravitational interactions with the central stellar component tend to modify the inner density profile [38, 39, 40].
Gondolo-Silk. Below the scale of tens of parsecs, DM is expected to form dense spikes. In their pivotal study [41], Gondolo & Silk (GS) showed that, assuming adiabatic accretion of the BH with DM particles gravitating in circular orbits around the central BH, an initial DM density profile of the form $\rho_{{\mathrm{DM}}}\propto r^{-\gamma}$ is modified into the steeper profile

### v3

@@ Abstract
Dual-phase xenon time projection chambers achieve optimal sensitivity for dark matter in the 10 to $1000\text{\,}\mathrm{G}\mathrm{eV}\mathrm{c}^{-2}$ mass range, but sub-GeV dark matter (DM) particles lack sufficient energy to produce nuclear recoils above detection thresholds in these detectors. Blazar-boosted dark matter offers a way to overcome this limitation. Relativistic jets in active galactic nuclei can accelerate light dark matter in their host-galaxy halos to energies that can leave detectable nuclear recoil signals in xenon-based detectors on Earth. We present the first blazar-boosted dark matter search that incorporates detector response modeling, using public data from XENON1T and LZ for the blazar TXS 0506+056. We model dark matter-proton scattering in the jet environment, covering the full process from jet acceleration through to detector response. We explore how the host galaxy dark matter density profile impacts our analysis. We set model-dependent exclusion regions on the dark-matter-nucleon scattering cross section for $m_{\chi}\simeq$1\text{\,}\mathrm{MeV}$\$ dark matter, between $5.8\text{\times}{10}^{-31}\text{\,}{\mathrm{cm}}^{2}$ and $6.3\text{\times}{10}^{-29}\text{\,}{\mathrm{cm}}^{2}$ using XENON1T data, and between $9.9\text{\times}{10}^{-32}\text{\,}{\mathrm{cm}}^{2}$ and $2.5\text{\times}{10}^{-28}\text{\,}{\mathrm{cm}}^{2}$ from LZ effective field theory (EFT) dark matter searches. Our results show that astrophysical uncertainties, especially those in the dark-matter distribution near the supermassive black hole, are the main limitation of this search rather than detector effects. The limits are therefore model-dependent and should be seen as exploratory. They highlight both the potential and the present uncertainties of blazar-boosted dark matter as a probe of light dark matter.
@@ I Introduction
The dark matter (DM) problem emerged from astronomical and cosmological observations that reveal a discrepancy between the observed gravitational effects and the visible matter in the Universe. This discrepancy requires either additional undetected mass or modifications to our current understanding of gravity. One proposed explanation involves dark matter particles, a form of nonbaryonic matter [Bertone2005Jan].
@@
Active galactic nuclei (AGNs) are the luminous central regions of galaxies powered by supermassive black holes. Blazars represent a special subclass of AGNs where jets are oriented toward Earth, making them appear exceptionally bright [Urry1995Sep, AGNReview].
@@
This paper is organized as follows: in Section II, we present the selection criteria for the blazar as a boosted-DM source and model both its jet physics and the DM profile of the host galaxy. In Sec. III, we calculate the BBDM flux, including the DM upscattering mechanism within the jet and Earth attenuation effects. Section IV describes the nuclear recoil interactions in liquid xenon detectors, essentially computing the recoil spectra. In Sec. V, we detail our approach for setting constraints on blazar-boosted dark matter using publicly available data from the XENON1T and LZ detectors. Section VI presents our conclusions.
@@ II.1 Selection of blazar-boosting source
In 2013, the IceCube Neutrino Observatory [icecubewebsite] detected the first high-energy astrophysical neutrino with energies exceeding $1\text{\,}\mathrm{P}\mathrm{e}\mathrm{V}$ [IceCube_2013]. Since then, it has compiled a growing list of such events (see Ref. [Icecat1] and references therein). Yet, so far, only the Galactic plane [icecube2023observation] (at 4.5 $\sigma$), and two astronomical objects have been associated with high-energy neutrinos detected by IceCube with a significance greater than $\sim 3\,\sigma$: the blazar TXS 0506+056  at $z=0.3365$ (at the 3–3.5 $\sigma$ level  [IceCube1, IceCube2], although more recent observations are consistent with a [not significantly] higher p-value [IceCube_ICRC_2024]) and the local Seyfert 2 galaxy NGC 1068 at $z=0.004$ (at the $4.2\,\sigma$ level  [IceCube:2022der, padovani2024high]). Many more lower-level associations between blazars and IceCube neutrinos have been suggested in the literature (e.g., [GP_2021, IceCube_2021, Sahakyan_2023] and references therein), which is consistent with a scenario where the contribution of the whole blazar class to the IceCube signal is not dominant but still relevant (e.g., Ref. [Bellenghi_2023]).
@@
Since light travel time from this blazar is 3.75 Gyr [1802.01939], any particles it emits must travel at least this long to reach Earth. Given that the Universe is 13.8 Gyr old, and that galaxies might form 0.2 Gyr after the big bang [e.g., Maiolino2024Mar], the blazar’s maximum possible age is about 9.8 Gyr (13.83.750.2=9.813.8-3.75-0.2=9.8 Gyr). For BBDM particles to be detectable on Earth, they must complete their journey within this 9.8 Gyr time frame. This constraint establishes a minimum velocity requirement for these particles, corresponding to a Lorentz factor of at least 1.1. Since this is very close to $\gamma=1$ (i.e., the nonrelativistic limit), this condition effectively places no meaningful constraint on the particles’ velocities.
@@ II.2 Blazar jet physics
The redshift of the source, and consequently its distance, has been firmly determined by Ref. [1802.01939].
@@
The mass of the central black hole (BH) in TXS 0506+056  is estimated to be $M_{\mathrm{BH}}\approx 6.3\times 10^{8}M_{\odot}$, where $M_{\odot}$ denotes the solar mass. This estimate is based on the assumption that the host galaxy is a typical giant elliptical with an absolute R magnitude of $M(R)\approx-22.9$ [Padovani_2022]. Consequently, this corresponds to an Eddington power $L_{\mathrm{Edd}}\approx 7.9\times 10^{46}$ erg s-1.
@@
The motion of the relativistic particles in the blazar jet can be approximately described according to the “blob geometry” [Dermer:2009zz] in which the accelerated particles are confined within spherical regions (blobs) that move with speed $v_{{\mathrm{B}}}=\beta_{{\mathrm{B}}}c$, and bulk Lorentz factor $\Gamma_{{\mathrm{B}}}=(1-\beta_{{\mathrm{B}}}^{2})^{-1/2}$ along the jet axis. The latter is, in general, inclined at an angle $\theta_{\text{l.o.s.}}$ with respect to our line-of-sight (l.o.s.). Within each blob, the individual particles move isotropically relative to the blob center of mass frame.
@@
In our analysis, we adopt the leptohadronic model used in [TXS, TXS_2] to describe the jet emission of blazar TXS 0506+056. This model characterizes the electromagnetic energy spectrum, i.e., the energy of the photons emitted by the accelerated charged particles (primarily electrons and protons) in the blazar’s jet. The model assumes a homogeneous and isotropic proton energy spectrum, following a single power-law with slope $\alpha_{p}$ within the energy range $\gamma^{\prime}_{{\mathrm{min}},p}\leq E_{p}^{\prime}/m_{p}\leq\gamma^{\prime}_{{\mathrm{max}},p}$, where $E^{\prime}_{p}$ is the proton energy in the blob’s rest frame. By fitting the observed electromagnetic spectral energy distribution, Ref. [TXS, TXS_2] determines the bulk Lorentz factor $\Gamma_{\rm B}$, the spectral index $\alpha_{p}$, the minimum and maximum Lorentz factors $\gamma^{\prime}_{{\mathrm{min}},p}$ and $\gamma^{\prime}_{{\mathrm{max}},p}$, and the proton luminosity $L_{p}$ that normalizes the energy power-law spectrum.
@@
The angle $\theta_{\text{l.o.s.}}$ for blazars is generally known to be small, with $\theta_{\text{l.o.s.}}<15-20^{\circ}$ (e.g. Ref. [UP_1995, AGNReview]). However, precise values for individual objects are uncertain and difficult to derive. In the analysis of [TXS, TXS_2], it is assumed that the bulk Lorentz factor of the jet $\Gamma_{\rm B}$ equals $\mathcal{D}/2$, where the latter is the Doppler factor $\mathcal{D}\equiv[\Gamma_{\rm B}\left(1-\beta_{\rm B}\cos\theta_{\text{l.o.s.}}\right)]^{-1}$. This implies $\theta_{\text{l.o.s.}}=0^{\circ}$, which is highly improbable and would preclude superluminal motion in this source [Padovani_2022_2]. This contradicts observations, as superluminal motion with $\beta_{\rm app}=1.07\pm 0.14$ has been measured [Lister_2021]. A further complication is that the jet parameters are very likely frequency-dependent, with $\Gamma_{\rm B}$ and $\mathcal{D}$ being much larger at high energies compared to the radio band (e.g. [Padovani_2022_2]).
@@
We therefore assume $\mathcal{D}=35$, which is the smallest value for this parameter derived by the lepto-hadronic spectral energy distribution modeling of [TXS_2], as this implies the largest possible $\theta_{\text{l.o.s.}}=\arcsin(1/\mathcal{D})\sim 1.64^{\circ}$ (see Appendix A of [UP_1995]). For this choice, and based on the definition $\mathcal{D}$, it can be shown that $\Gamma_{\rm B}=\mathcal{D}$. This means that the lowest value for the proton power in Table 1 of Ref. [TXS_2], which has been derived assuming $\Gamma_{\rm B}=\mathcal{D}/2$, needs to be multiplied by a factor of 4 for consistency, giving $L_{p}=6.4\times 10^{48}\,\text{erg}/\text{s}$. We take the other blazar jet parameters at the best-fit values given in [TXS_2]. The entire list of parameters and corresponding values considered in our work are summarized in Table 1.
@@
With the parameters in the table, we estimate the number of protons ejected per time, kinetic energy $T_{p}$ and solid angle $\Omega_{p}$, in the rest frame of an observer that sees the blobs moving with velocity $\beta_{{\mathrm{B}}}$ along the jet axis, as [Wang:2021jic]
@@
where $m_{p}\simeq 0.938\,\text{GeV}/c^{2}$ is the proton mass, $\beta_{p}=[1-m^{2}_{p}/(T_{p}+m_{p})^{2}]^{1/2}$ is the proton speed in units of $c$, $\mu$ is the cosine of the angle of motion with respect to the jet axis, and $c_{p}$ is a normalization constant, which we fix by requiring the matching with the fitted proton luminosity $L_{p}$ [Granelli:2022ysi].
@@
Table 1: Parameters currently used in the BBDM code for TXS 0506+056 [TXS, TXS_2]. These parameters were partially derived from [Wang:2021jic, Granelli:2022ysi]. The quantities flagged with a star () correspond to mean values computed from the ranges given in the second column of Table 1 of Ref. [TXS_2]. The resulting values of the normalization constants $c_{e,\,p}$, as well as the redshift $z$ [Keivani:2018rnh], luminosity distance $d_{L}$ for $H_{0}=70$, $\Omega_{\rm m}=0.3$, and $\Omega_{\rm\Lambda}=0.7$, BH mass $M_{{\mathrm{BH}}}$ [Padovani_2022].
@@ II.3 Dark matter density profile
Figure 1 shows the dark matter density and integrated column density for the three models.
II.3.1 Navarro-FrenkWhite
The NFW profile [Navarro:1996gj] is often used as the default model for galactic DM halos. It takes the form:
@@ II.3.1 Navarro-FrenkWhite
where $\rho_{s}$ is the characteristic density (a normalization parameter that sets the density scale) and $r_{s}$ is the characteristic scale radius (which determines the transition point between the inner and outer regions of the density profile). The characteristic density $\rho_{s}$ is defined [Gentile:2007sb] by requiring that the total mass within the virial radius $r_{\text{vir}}$ equals the virial mass:
@@
where $\delta_{\text{vir}}\simeq 100$ is the virial overdensity in the standard cosmological model [Bryan:1997dn], and $\rho_{c}=1.053672\times 10^{-5}\,h^{-2}\,(\text{GeV}/c^{2})\,\text{cm}^{-3}$ is the critical density of the Universe. Here, $h\simeq 0.674$ is the Hubble expansion rate in units of 100 km/s/Mpc [PDG_Workman:2022ynf].
@@
The characteristic radius $r_{s}$ and the virial radius $r_{\text{vir}}$ entering Eq.  (5) are expressed in terms of the virial mass as[Wechsler:2005gb, Klypin:2010qw]
@@
The virial mass of the DM halo is estimated via the following relation which stems from numerical simulations [DiMatteo:2003zx] (see also [Ferrarese:2002ct, Baes:2003rt]):
@@
giving $M_{\text{vir}}\simeq 5.2\times 10^{12}M_{\odot}$ for the parameters in Table 1. Using this value in Eqs. (7) and (6), we get $r_{\text{vir}}\simeq 273\,\text{kpc}$ and $r_{s}\simeq 54\,\text{kpc}$, while from Eq. (5), $\rho_{s}\simeq 0.10\,(\text{GeV}/c^{2})\,\text{cm}^{-3}$.
The strength of the NFW profile lies in its close agreement with observed rotation velocity curves at large radii. However, at small radii, the DM profile is poorly constrained, making the NFW profile an extrapolation of large-scale trends to small scales without strong empirical support. Furthermore, numerical simulations indicate that the NFW profile is not expected to hold at small radii, since baryonic processes such as stellar feedback and gravitational interactions with the central stellar component tend to modify the inner density profile [Gnedin2004Aug, Merritt2007Feb, Shapiro2022Aug].
II.3.2 Gondolo-Silk
Below the scale of tens of parsecs, DM is expected to form dense spikes. In their pivotal study [Gondolo:1999ef], Gondolo and Silk (GS) showed that, assuming adiabatic accretion of the BH with DM particles gravitating in circular orbits around the central BH, an initial DM density profile of the form $\rho_{{\mathrm{DM}}}\propto r^{-\gamma}$ is modified into the steeper profile